# Two distinct types of echoes in compact objects

### v1

@@ I Introduction
Cardoso et al. introduced the concept of echoes [38], a novel phenomenon intersecting with late-stage ringing waveforms, potentially distinguishing different gravitational systems through unique near-horizon characteristics. This concept has spurred numerous investigations into echoes across various systems, including exotic compact objects such as gravastars [39, 40], wormholes [41, 42, 43, 44, 45], among others.
@@ II Two physical pictures for echoes in compact objects
II.1 The first type of echoes
@@
II.2 The second type of echoes
Now, we turn our attention to the second type of echoes. In this case, the emergence of echoes is closely associated with the presence of a discontinuity in the effective potential. The discontinuity that triggers these echoes may be positioned near what would traditionally be considered the horizon or at a more distant location relative to the compact object. The core principles underlying the formulation of these echoes remain consistent across different physical systems. As discussed previously, the mathematical framework for deriving echoes from exotic compact objects has been established in [46, 48].
@@ II.2 The second type of echoes
The succeeding arguments are mainly reminiscent of the preceding subsection. Specifically, if $\omega$ is a pole of Eq. (21) and therefore the Green’s function Eq. (22), $\omega+n\pi/x_{c}$ (where $n$ is an arbitrary integer) is also a pole. The union of such poles gives rise to an additional branch of QNMs, contributing to echoes of a period of $2x_{c}$.
@@ V Concluding remarks
As implemented by this unified model, the investigation of echoes in compact stars enriches our theoretical toolkit for probing the universe’s most extreme gravitational environments. It potentially opens up new avenues for testing the predictions of general relativity and exploring alternative theories of gravity, thereby contributing to our ongoing quest to understand the cosmos.

### v2

@@ I Introduction
Cardoso et al. introduced the concept of echoes [38], a novel phenomenon intersecting with late-stage ringing waveforms, potentially distinguishing different gravitational systems through unique near-horizon characteristics. This concept has spurred numerous investigations [39, 40, 41] into echoes across various systems, including exotic compact objects such as gravastars [42, 43], wormholes [44, 45, 46, 47, 48, 49, 50], among others.
@@ II Two physical pictures for echoes in compact objects
II.1 The first type of echo
@@
II.2 The second type of echo
Now, we turn our attention to the second type of echo. In this case, the emergence of echoes is closely associated with the presence of a discontinuity in the effective potential. The discontinuity that triggers these echoes may be positioned near what would traditionally be considered the horizon or at a more distant location relative to the compact object. The core principles underlying the formulation of these echoes remain consistent across different physical systems. As discussed previously, the mathematical framework for deriving echoes from exotic compact objects has been established in [51, 53].
@@ II.2 The second type of echo
The succeeding arguments are mainly reminiscent of the preceding subsection. Specifically, if $\omega$ is a pole of Eq. (21) and therefore the Green’s function Eq. (22), $\omega+n\pi/x_{c}$ (where $n$ is an arbitrary integer) is also a pole. The union of such poles gives rise to an additional branch of QNMs, contributing to echoes of a period of $2x_{c}$ [^1] Rather than representing an absolute coordinate, we note that $x_{c}$ essentially measures a relative distance. It quantifies the offset between the location of the effective potential, relative to which the transmission coefficients are defined and evaluated, and the discontinuity. This has been implied in the preceding derivations..
@@ V Concluding remarks
For a given total mass, the results presented in Sec. IV indicate that the second type of echo becomes relevant only when the star’s size is significant. Specifically, the criterion for the emergence of the second type of echo is that the star’s surface lies outside the would-have-been local maximum of the Regge-Wheeler effective potential, i.e., $r_{b}\gtrapprox r_{\mathrm{max}}$, where $r_{\mathrm{max}}=3M$ in the eikonal limit. At the same time, one might pursue other observational channels. The above criterion ensures that such an object will likely not possess a photon sphere, specifically $r_{p}=3M$ for the Regge-Wheeler potential, resulting in a distinct optical image compared to objects emitting the first type of echo. From an empirical perspective, we argue that this condition is feasible in realistic scenarios, and various compact objects may exhibit the second type of echo. For instance, neutron stars typically have a radius of 10-15 kilometers, while their Schwarzschild radius is about 3 kilometers [81]. This implies that neutron stars usually do not have a photon sphere and are subject to the second type of echo. On the other hand, as the density of the compact object increases further, it might become optically “dark” due to the presence of a photon sphere while exhibiting the first type of echo. Furthermore, beyond the specific quasinormal oscillations of stars discussed in this work, we speculate that the second type of echo signal could have broader relevance in a universal context. For instance, on a larger scale, discontinuities arising from phenomena like Bondi accretion around compact objects could give rise to such echoes, which might be used to probe their presence. Similarly, these echoes might manifest in the gravitational wave spectrum of compact binaries, as surfaces and resulting discontinuities are inherent features of any compact object. In conclusion, our findings regarding the two types of echoes in compact stars enrich the theoretical framework for investigating the universe’s most extreme gravitational environments. This potentially opens new pathways for testing the predictions of general relativity and exploring alternative theories of gravity, contributing to our ongoing quest to understand the cosmos.